OpenStudy (gabylovesyou):

Point A is located at (0, 4) and point B is located at (−2, −3). Find the x value for the point that is 1/4 the distance from point A to point B. −1 −0.75 −0.5 −0.25

OpenStudy (gabylovesyou):

@johnweldon1993

OpenStudy (anonymous):

-0.25

OpenStudy (gabylovesyou):

Oh wow hi ^.^ how did u get that ?

OpenStudy (anonymous):

1/4=0.25 so in your equation it has to cancel out so you would subtract it (-0.25) i may be wrong but that is how my math teacher taught me

OpenStudy (gabylovesyou):

Thanks! well ill just make sure then ... @ParthKohli

OpenStudy (anonymous):

@mathmale

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (gabylovesyou):

yes ..

OpenStudy (gabylovesyou):

thank you!

OpenStudy (anonymous):

your absolutly not welcome lol

OpenStudy (johnweldon1993):

Ehh nvm...sooo incorrect I believe lol

OpenStudy (gabylovesyou):

hmmm lol

OpenStudy (anonymous):

^^^ its -0.25 right

OpenStudy (johnweldon1993):

Oh wait no...hahaha I read it right XD sorry I'm going back and forth

OpenStudy (gabylovesyou):

LMFAO @jzellem1004 HA xD

OpenStudy (anonymous):

mhm lol

OpenStudy (gabylovesyou):

i tagged someone else lol @johnweldon1993 thanks

OpenStudy (johnweldon1993):

Lol I was reading it wrong the second time.... So yeah it is just finding the distance between the points...and 1/4 that in regards to the 'x' value...sorry @Gabylovesyou :) But yes -0.5 is correct :)

OpenStudy (anonymous):

tell her that she isnt welcome

OpenStudy (gabylovesyou):

so if i wanted to answer this one, i put a 2 or is this asking something else ? Find the x value for the point that splits segment AB in half if point A is located at (4, 3) and point B is located at (−2, 6).

OpenStudy (johnweldon1993):

@Gabylovesyou you are VERY welcome :)

OpenStudy (gabylovesyou):

@jzellem1004 shush lol

OpenStudy (johnweldon1993):

Right...because that is just finding the midpoint of said segment

OpenStudy (anonymous):

nope @johnweldon1993 she isnt welcome ok ur being pushed around @Gabylovesyou he messaged me saying u werent welcome soz

OpenStudy (johnweldon1993):

\[\large \frac{x_2 + x_1}{2} = \text{Midpoint}\]

OpenStudy (anonymous):

and midpoint= you aint welcome

OpenStudy (gabylovesyou):

lol ok ummm so it would be 4 + -2 / 2 ?

OpenStudy (anonymous):

which is =1

OpenStudy (johnweldon1993):

Correct @Gabylovesyou which evaluates out to...?

OpenStudy (gabylovesyou):

4 + -2 = 2 ? which is 2/2 = 1 ?

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (anonymous):

i already said that

OpenStudy (gabylovesyou):

yaay thanks

OpenStudy (anonymous):

NOT PERFECT 10 IS PERFECT U SAID ONE lololol

OpenStudy (the_fizicx99):

o_o ...

OpenStudy (gabylovesyou):

wut lol @jzellem1004

OpenStudy (anonymous):

on sw when u were rating u said 10 is perfect and no one is perfect lol

OpenStudy (gabylovesyou):

oh lmfao xD

OpenStudy (anonymous):

guess what @Gabylovesyou

OpenStudy (gabylovesyou):

what @jzellem1004

OpenStudy (anonymous):

ur pretty :P

OpenStudy (gabylovesyou):

lmfao xD PC me -.-

OpenStudy (anonymous):

WELL UR NOT WELCOME, o u didnt say thanks v.v